OpenStudy (anonymous):

3. Find two numbers whose sum is 18 such that when the larger is divided by the smaller, the quotient is 2 and the remainder is 3.

OpenStudy (anonymous):

x + y = 18, x-3/y =2, 2y = x-3, x = 2y + 3, substitute x into x + y = 18 u get 2y + 3 + y = 18.

OpenStudy (anonymous):

3y = 15, y = 5, x is 13

OpenStudy (anonymous):

thank you so much :D